# Pcodec: Better Compression for Numerical Sequences

### v1

@@ Abstract
We present Pcodec (Pco), a format and algorithm for losslessly compressing numerical sequences. Pco’s core and most novel component is a binning algorithm that quickly converges to the true entropy of smoothly, independently, and identically distributed (SIID) data. To automatically handle more general data, Pco has two opinionated preprocessing steps. The first step, Pco’s mode, decomposes the data into more smoothly distributed latent variables. The second step, delta encoding, makes the latents more independently and identically distributed. We prove that, given $k$ bins, binning uses only $\mathcal{O}(1/k)$ bits more than the SIID data’s entropy. Additionally, we demonstrate that Pco achieves 29-94% higher compression ratio than other approaches on six real-world datasets while using less compression time.
@@ 1 Introduction
Enormous volumes of numerical data are stored in columnar, tensor, and time series formats, with even individual datasets measured in petabytes and exabytes [1, 2, 3]. Compression improvements offer tremendous value in freed data storage. And since these datasets are often consumed by query engines and machine learning workloads from disk or network devices, better compression can improve performance as well [4].
@@
Much prior research on lossless numerical compression has focused on compression and decompression speed rather than compression ratio [5, 6, 7, 8]. These approaches are particularly suitable for data transfers between CPU cores, since per-core memory bandwidth is high. However, when reading from disk or network devices, per-core bandwidth is often under 0.1 GiB/s. In this regime, CPU can cease to be a bottleneck, leaving compression ratio as the only performance consideration. Additionally, when compressing for storage, costs are directly proportional to file size, making compression ratio paramount.
@@
Relatively little research has been done to address this need for high compression ratios on numerical sequences. All such approaches we know of use general-purpose LZ77 [9] or LZ78 [10] (hereafter referred to as LZ) codecs as a final step. Popular columnar formats such as Parquet [11] support a single preprocessing step prior to LZ compression. Examples of Parquet’s preprocessing steps include Dictionary, which encodes each value as an index into a deduplicated list of unique values, and Delta, which takes differences between nearby numbers and packs them into variable-length integers. Similar to Parquet Delta, the Turbo PFor library offers its own bit packing algorithms [12]. Blosc2, a tensor format, supports preprocessing filters such as Shuffle [^1] Parquet has a similar encoding known as ByteStreamSplit, but does not support it for integer data types., which transposes the data such that all numbers’ first bytes come first, followed by their second bytes, etc [13].
@@
These preprocessing steps all commonly pair with industry-standard LZ codecs such as Zstandard (Zstd) [14]. Another possible LZ codec is SPDP, whose parameters were tuned to suit numerical data [15]. However, all LZ codecs fundamentally treat bytes as discrete symbols, discarding numerical information. For instance, they have no a priori understanding that the byte 3 is closer to 4 than to 255. LZ codecs are also unaware of the boundaries between numbers. This suggests that there may be room for improvement.
@@
We present Pcodec (Pco), a lossless format and algorithm designed to better compress numerical sequences. The core idea of Pco is to express the input numbers in terms of smoothly (defined in Section 3.1), independently, and identically distributed (SIID) unsigned integers. We call these unsigned integers latents. Pco compresses these latents with a technique called binning, which quickly converges to the true entropy of any SIID data. Contrast this with the core idea of most LZ codecs, which is to express the input bytes in terms of independently and identically distributed (IID) symbols from a small alphabet, and to use entropy coding techniques to converge to the Shannon entropy of the symbol distribution. Intuitively, we expect Pco to outperform LZ codecs on numerical distributions because expressing numbers as SIID latents is more natural than expressing them as IID symbols.
Pco does this via two preprocessing steps: its mode and delta encoding. By default, Pco chooses good configurations for these preprocessing steps automatically, using robust algorithms. These preprocessing steps are opinionated: each chunk of data has exactly one mode and one delta encoding.
@@
On a wide range of real-world datasets, we find that Pco obtains 29% to 94% higher compression ratio than alternatives, even when granting those alternatives 50% more compression time. Additionally, Pco’s decompression speeds consistently surpassed 1GiB/s per thread.
@@
Early work on Pco started in May 2021. Today, it is used in tools like Zarr [16] and CnosDB [17] to store petabytes of data. Source code is available at https://github.com/mwlon/pcodec/.
@@ 2.1 Modes
At present, Pco has four modes: Classic, IntMult, FloatMult, and FloatQuant (Table 1). For brevity, we describe only Classic and IntMult in depth, but the others are similar in principle.
Mode Data Types Example Use Case
@@
IntMult integers millisecond-precise timestamps stored as microseconds
FloatMult floats prices that are mostly multiples of 0.01
@@
2.1.1 Splitting and Joining
Each mode’s purpose is to split the distribution of numbers into one or two latent variables during compression, then join them back into numbers during decompression. These splitting and joining operations form a bijection. Although splitting the data into two latent variables can only increase its theoretical entropy, it can reduce the number of bits used by binning.
@@ 2.0 $010\ldots 0$ $110\ldots 0$
Table 2: Examples of how Classic mode preserves order when converting floats to latents. The IEEE representation of each float is not monotonic in the float’s numerical value, but the latent representation is.
Classic mode is tailored for numbers that are already smoothly distributed. Therefore, it simply splits each number to a single latent in an order-preserving way. Since latents are always unsigned integers, this can actually be a nontrivial function. For instance, a positive float is converted to a latent by toggling its sign bit, while a negative float has all bits toggled (Table 2).
IntMult mode handles integer distributions that resemble a multiply-add of two latent variables, i.e. $x_{i}=q_{i}m+r_{i}$ for some parameter $m$ and smooth latent distributions $q_{i}\sim Q_{m}$, $r_{i}\sim R_{m}$. For instance, if 90% of the numbers are congruent to 7 modulo 101, we could benefit from IntMult mode with $m=101$, splitting the data into latents $q_{i}=\lfloor x_{i}/101\rfloor$ and $r_{i}$ is $x_{i}$’s remainder mod 101.
@@ 2.1.2 Automatic Detection
When the user does not configure the mode, Pco chooses it via statistical techniques on a deterministic random sample containing a few percent of the numbers. Each non-Classic mode chooses a parameter and an estimate of how many bits it would save over Classic. Pco proceeds with the mode with the lowest estimated bit cost.
Consider IntMult mode. To propose values for its parameter $m$, we group our sample of $n_{\text{sample}}$ numbers into $c=\lfloor n_{\text{sample}}/3\rfloor$ triples, then consider $m=\gcd(x_{2}-x_{1},x_{3}-x_{1})$ for each triple $(x_{1},x_{2},x_{3})$. Values of $m$ that occur anomalously frequently make good candidates. Let $c_{m}$ be the number of times the GCD $m$ appeared, and let $n_{\text{infrequent}}$ be the count of numbers $x_{i}$ from the sample with at most $n_{\text{sample}}/256$ sharing their value for $q_{i}$. Using statistical techniques (Appendix B), we obtain conservative approximations for the bit costs of the quotient distribution $Q_{m}$ and remainder distribution $R_{m}$:
@@ 2.2 Delta Encoding
After applying the mode, Pco uses delta encoding to disentangle correlations between numbers, making them more IID. At present, Pco has 3 delta encodings:
@@
  • Lookback: Chooses a lookback $l_{i}$ for each latent, then transforms each latent to be $y^{\prime}_{i}=y_{i}-y_{i-l_{i}}$. The lookbacks are stored as an additional latent variable, and therefore get binned as well. Lookback allows Pco to capture LZ-like patterns at an approximate number level, as opposed to at an exact byte level.
@@ 2.2.1 Automatic Detection
When the user does not configure delta encoding, Pco chooses it by measuring compressed size of different configurations on a sample of data. This sample is composed of multiple runs of roughly 100 consecutive numbers each. Because Consecutive with $o>1$ is rarely advantageous, we try orders $o=1,\ldots,7$ sequentially and stop early as soon as the compressed size with $o$ exceeds that of $o-1$. Once it has computed the sample’s compressed size for each configuration, Pco simply chooses the delta encoding with the lowest size.
@@ 2.3 Binning
Pco’s most essential algorithm is binning, which compresses SIID latent variables. It is the only processing step in Pco that reduces the byte size of the data. The concept is to choose histogram bins over the latent variable, then encode each latent by its entropy-coded bin and offset into that bin.
Binning takes as input a vector of latents $(y_{1},\ldots,y_{n})$. The first step is to compute a histogram of $k$ tight, exclusive bins $(a_{1},b_{1},c_{1}),\ldots,(a_{k},b_{k},c_{k})$, where $a_{i}$ is a lower bound, $b_{i}$ is an upper bound, and $c_{i}$ is the count of latents in the bin. By tight we mean that for every $i$ there exist some $y_{j}$ and $y_{l}$ such that $y_{j}=a_{i},y_{l}=b_{i}$, and by exclusive we mean that $b_{i}<a_{i+1}$.
The second step is to optimize these bins, resulting in $(a^{\prime}_{1},b^{\prime}_{1},c^{\prime}_{1}),\ldots,(a^{\prime}_{k^{\prime}},b^{\prime}_{k^{\prime}},c^{\prime}_{k^{\prime}})$, which we explain in Section 2.3.1. We build a tANS table [18] to entropy code the optimized bins, with bin $j$ weighted by $c^{\prime}_{j}$.
Finally, we write the $y$ values. ANS codes must be decoded in the opposite order they are computed, so we first iterate through each $y_{i}$ in reverse, find the $j$ such that $a^{\prime}_{j}\leq y_{i}\leq b^{\prime}_{j}$, and compute a tANS code for it. Then we iterate through the $y$ values forwards in batches of 256. For each $y_{i}$ we write out the previously-computed tANS code and record the offset $y_{i}-a^{\prime}_{j}$ as a $\lceil\log_{2}(b^{\prime}_{j}-a^{\prime}_{j}+1)\rceil$-bit unsigned integer. Once all tANS codes in a batch are written, we write these offsets.
@@ 2.3.1 Bin Optimization
Since each additional bin incurs some metadata storage, it is usually suboptimal to store all histogram bins. Adjacent bins often have very similar probability density, so they can be combined.
To combine and optimize the bins, we use an $\mathcal{O}(k^{2})$ dynamic programming algorithm [19] to partition the bins and minimize the final compressed bit size, taking metadata, tANS bits, and offset bits into account. This algorithm requires only a cost function – in our case, the total bit cost of all numbers to be stored in an optimized bin. The cost of the optimized bin formed by combining unoptimized bins $i$ through $j$ is
$\hat{H}_{i:j}=M+c_{i:j}\cdot(\alpha_{i:j}+\beta_{i:j})$
where $M$ is a constant for the number of bits required to store the bin’s metadata, $c_{i:j}=\sum_{l=i}^{j}c_{l}$ is the count of latents in the bin, $\alpha_{i:j}=\log_{2}\left(c/c_{i:j}\right)$ is the idealized average entropy coding bits per latent in the bin, and $\beta_{i:j}=\lceil\log_{2}(b_{j}-a_{i}+1)\rceil$ is the offset bits per latent in the bin. Essentially, the bin pays its metadata cost and, for each number in its range, entropy coding and offset cost.
Once the optimal bin partitioning is chosen, we replace each partition’s bins $(a_{i},b_{i},c_{i}),\ldots,(a_{j},b_{j},c_{j})$ with a single bin
$(a^{\prime},b^{\prime},c^{\prime})=\left(a_{i},b_{j},\sum_{l=i}^{j}c_{l}\right)$
@@ 2.4 Format
Pco is most commonly used within a wrapping format. We found it important to decouple Pco’s goal of good compression from the varied access patterns and logical transformations required by real-world applications, which may include nullability, page statistics, or additional indices. To support these needs, Pco’s API produces three components that can be interleaved as needed.
@@
Every chunk contains one or more pages, which contain the actual numerical data (Figure 2). Pages store initial states for tANS and delta decoding, so when paired with the correct chunk metadata, they contain all information necessary to decompress their contents. Therefore pages can be decompressed in any order; it is not necessary to decompress preceding pages.
Each page can contain multiple batches, the unit of decompression. Batches consist of all latents for 256 numbers and exist solely for performance purposes. Batches within a page must be decompressed serially.
@@ 3.1 Theoretical Results
In Appendix A, we prove a bound on Pco’s bit cost per number:
Theorem 1
Suppose $X$ is mixture over $\{0,\ldots,T-1\}$ of $s$ disjoint integer distributions, each of which has a monotonic PMF. Then for any $k>2s$, there exists a binning of at most $k$ bins such that the binned bit cost $\hat{H}$ of a random draw from $X$ satisfies
$\hat{H}\leq H+\frac{5s\log_{2}(T)}{k-2s}$
where $H$ is the base-2 Shannon entropy of $X$.
We have proved this by construction, and the binning that attains this bound uses bins of approximately equal probability mass. This lines up with our approach of using a histogram to assign unoptimized bins. Informally, we call $X$ smooth if $s$ is small. And this is the case for many types of data; geometric distributions have $s=1$, Poisson distributions have $s=2$, and normal distributions have $s\in\{2,4\}$.
Pco converges faster than this bound in reality (Figure 3). However, this bound can still be practical. For instance, using the default compression level’s 256 bins on a 64-bit data type with $s=1$, we guarantee that binning wastes at most $160/127\approx 1.26$ bits per latent on average.

### v2

@@ Abstract
We present Pcodec (Pco), a format and algorithm for losslessly compressing numerical (float or integer) sequences. Pco’s core and most novel component is a binning algorithm that quickly converges to the true entropy of smoothly, independently, and identically distributed (SIID) integers. We mathematically prove this convergence with a practical bound. To accommodate data this is not SIID, Pco has two opinionated preprocessing steps. The first step, Pco’s mode, decomposes the numbers into more smoothly distributed integer latent variables. The second step, delta encoding, makes the latents more independently and identically distributed. We demonstrate that Pco achieves 29-94% higher compression ratio than other numerical codecs on six real-world columnar datasets while using less compression time.
@@ 1 Introduction
Overwhelming quantities of numerical data are stored in columnar, tensor, and time series formats, with even individual datasets measured in petabytes and exabytes [1, 2, 3]. Compression improvements offer tremendous value in freed data storage. And since these datasets are often consumed by query engines and machine learning workloads from disk or network devices, better compression can improve performance as well [4].
@@
Much prior research on lossless numerical compression has focused on speed rather than compression ratio [5, 6, 7, 8]. These approaches are particularly suitable for data transfers between CPU cores, where per-core memory bandwidth is high. However, when reading from disk or network devices, the optimal balance may require stronger compression. In this regime, CPU can cease to be a bottleneck, leaving compression ratio as the only performance consideration. Additionally, compression ratio is paramount when compressing for long-term storage, where costs are directly proportional to file size. Relatively little research has been done to address this need for high compression ratios on general numerical sequences. Most such approaches we know of use general-purpose LZ77 [9] or LZ78 [10] (hereafter referred to as LZ) codecs as a final step. We will briefly review several popular or published codecs in this space.
@@
The columnar format Parquet [11] supports a preprocessing step prior to LZ compression. Examples of Parquet’s encodings include Dictionary, which encodes each number as an index into a deduplicated list of unique values, and Delta, which takes differences between nearby numbers (a type of delta encoding) and packs them into variable-length integers. BtrBlocks built on this idea by adding more encodings and using small samples of data to automatically choose a good encoding [12]. Uniquely among the codecs listed here, BtrBlocks eshews LZ compression. The Turbo PFor library offers a variety of preprocessing steps as well, including its namesake Patched Frame of reference algorithm [13]. PFor algorithms store small numbers using a fixed bit width and store large exceptions separately. Blosc2, a tensor format, supports the Shuffle [^1] Parquet has a similar encoding known as ByteStreamSplit, but does not support it for integer data types. preprocessing filter, which transposes the data such that all numbers’ first bytes come first, followed by their second bytes, etc [14].
@@
Two LZ codecs are worth mentioning. First, Zstandard (Zstd) [15] is an industry-standard LZ codec, used either by itself or paired with one of the aforementioned preprocessing steps. SPDP is another LZ codec whose parameters were tuned to suit numerical data, and therefore should be used by itself [16]. However, all LZ codecs fundamentally treat bytes as discrete symbols, discarding numerical information. They have no a priori understanding that the byte 3 is closer to 4 than to 255. LZ codecs are also unaware of the boundaries between numbers. This suggests that there may be room for improvement.
@@
We present Pcodec (Pco), a lossless format and algorithm designed to better compress numerical sequences. The core idea of Pco is to express the input numbers in terms of smoothly (defined in Section 3.1), independently, and identically distributed (SIID) unsigned integers. We call these unsigned integers latents. Pco compresses latents with a technique called binning, which quickly converges to the true entropy of any SIID data (Section 3.1). Contrast this with the core idea of most LZ codecs, which is to express repetitions of the input bytes in terms of independently and identically distributed (IID) symbols from a small alphabet, and to use entropy coding techniques to converge to the Shannon entropy of the symbol distribution. Intuitively, we expect Pco to outperform LZ codecs on numerical distributions because it is more natural to express numbers in terms of mathematical formulas on latents than in terms of repeated byte sequences. Pco decomposes numbers into latents via two preprocessing steps: its mode and delta encoding. By default, Pco uses robust algorithms to automatically choose good configurations for these. This process is opinionated; each chunk of data can use exactly one mode and one delta encoding.
@@
On a wide range of real-world datasets, we find that Pco obtains 29% to 94% higher compression ratio than alternatives, even when granting those alternatives 50% more compression time. Additionally, Pco’s decompression speeds consistently surpass 1GiB/s per thread.
@@
Early work on Pco started in May 2021. Today, it is used in tools like Zarr [17] and CnosDB [18] to store petabytes of data. Source code is available at https://github.com/pcodec/pcodec, including a CLI with benchmarks for all codecs we compare against.
@@ 2.1 Modes
At present, Pco has four modes: Classic, IntMult, FloatMult, and FloatQuant (Table 1). Each mode’s purpose is to exploit common structure in the numbers by splitting the distribution into one or two latent variables during compression. During decompression, the mode joins these latents back into the original numbers. Although the sum of entropy over latent variables must be at least as high as the original distribution’s entropy, the latent variables can be smoother, reducing the number of bits used downstream during binning.
Mode Types Example Use Case
@@
IntMult integers ms-precise timestamps stored as µs
FloatMult floats prices that are multiples of 0.01
@@
For brevity, we describe only Classic and IntMult in depth, but the others are similar in principle. Classic mode is tailored for numbers that are already smoothly distributed. Therefore, it simply splits each number to a single latent in an order-preserving way. For instance, a positive float is converted to a latent by toggling its sign bit, while a negative float has all bits toggled (Table 2).
@@ 2.0 $010\ldots 0$ $110\ldots 0$
Table 2: Examples of how Classic mode preserves order when converting floats to latents. The IEEE representation of each float is non-monotonic in the float’s numerical value, but the latent representation is.
IntMult mode handles integer distributions that resemble a multiply-add of two latent variables, i.e. $x_{i}=q_{i}m+r_{i}$ for some parameter $m$ and latent distributions $q_{i}\sim Q_{m}$, $r_{i}\sim R_{m}$. For instance, it could be beneficial for a jagged distribution where 90% of numbers are congruent to 7 modulo 101. By using $m=101$, IntMult mode would let $q_{i}$ be $\lfloor x_{i}/101\rfloor$ and $r_{i}$ be $x_{i}$’s remainder mod 101. The resulting $Q_{m}$ and $R_{m}$ distributions would be smoother than the original distribution.
@@ 2.1.1 Automatic Detection
When the user does not configure the mode, Pco chooses it via statistical techniques on a deterministic random sample containing a few percent of the numbers. Each non-Classic mode chooses a parameter and an estimate of how many bits it would save over Classic. Pco proceeds with the mode with the lowest estimated number of bits required, which we refer to as the bit cost, denoted as $\hat{H}$.
Consider IntMult mode. To propose values for its parameter $m$, we group our sample of $n_{\text{sample}}$ numbers into $c=\lfloor n_{\text{sample}}/3\rfloor$ triples, then let $m=\gcd(x_{2}-x_{1},x_{3}-x_{1})$ for each triple $(x_{1},x_{2},x_{3})$. It follows that $x_{2}$ is congruent to $x_{3}$ modulo $m$, so if some value of $m$ appears for anomalously many triples, it suggests that the quotient distribution $Q_{m}$ might be low-entropy. With that intuition, we search for frequent values of $m$ among the triples. In Appendix B, we derive conservative approximations for the bit costs of $Q_{m}$ and $R_{m}$, leveraging [19]. Let $c_{m}$ be the number of times the GCD $m$ appeared, and let $n_{\text{infrequent}}$ be the count of numbers $x_{i}$ from the sample with at most $n_{\text{sample}}/256$ sharing their value for $q_{i}$. Then the bit cost of $Q_{m}$ can be compared to that of Classic mode:
@@ 2.2 Delta Encoding
After applying the mode, Pco uses delta encoding to disentangle correlations between latents within each variable, making them more IID. At present, Pco has 3 delta encodings:
@@
  • Lookback: Chooses a lookback $l_{i}$ for each latent, then transforms each latent $y_{i}$ to be $y^{\prime}_{i}=y_{i}-y_{i-l_{i}}$. The lookbacks are stored as an additional latent variable, and therefore get binned as well. Lookback allows Pco to capture LZ-like patterns at an approximate number level, as opposed to at an exact byte level.
@@ 2.2.1 Automatic Detection
When the user does not configure delta encoding, Pco chooses it by measuring compressed size of different configurations on a sample of data. In order to capture correlations between nearby latents, this sample is composed of multiple runs of roughly 100 consecutive numbers each. Because Consecutive with $o>1$ is rarely advantageous, we try orders $o=1,\ldots,7$ sequentially and stop early as soon as the compressed size with $o$ exceeds that of $o-1$. Once it has computed the sample’s compressed size for each configuration, Pco simply chooses the delta encoding with the lowest size.
@@ 2.3 Binning
Pco’s most essential algorithm is binning, which compresses SIID latent variables. It is the only processing step in Pco that reduces the byte size of the data. The overarching idea is to choose histogram bins over the latent variable, then encode each latent by its entropy-coded bin and offset into that bin.
Binning takes as input a vector of latents $(y_{1},\ldots,y_{n})$. The first step is to compute a histogram of $k$ tight, exclusive bins $(a_{1},b_{1},n_{1}),\allowbreak\ldots,\allowbreak(a_{k},b_{k},n_{k})$, where $a_{i}$ is a lower bound, $b_{i}$ is an upper bound, and $n_{i}$ is the count of latents in the bin. By tight we mean that for every $i$ there exist some $y_{j}$ and $y_{l}$ such that $y_{j}=a_{i},y_{l}=b_{i}$, and by exclusive we mean that $b_{i}<a_{i+1}$.
The second step is to optimize these bins, resulting in $(a^{\prime}_{1},b^{\prime}_{1},n^{\prime}_{1}),\allowbreak\ldots,\allowbreak(a^{\prime}_{k^{\prime}},b^{\prime}_{k^{\prime}},n^{\prime}_{k^{\prime}})$, which we explain in Section 2.3.1. We build a tANS table [20] to entropy code the optimized bins, with bin $j$ weighted by $n^{\prime}_{j}$.
Finally, we write the $y$ values. ANS codes must be decoded in the opposite order they are encoded, so we encode in reverse. For each $y_{i}$, we find the $j$ such that $a^{\prime}_{j}\leq y_{i}\leq b^{\prime}_{j}$ and compute a tANS code for it. Then we iterate through the $y$ values forwards in batches of 256. For each $y_{i}$ we write out the previously-computed tANS code and record the offset $y_{i}-a^{\prime}_{j}$ as a $\lceil\log_{2}(b^{\prime}_{j}-a^{\prime}_{j}+1)\rceil$-bit unsigned integer. Once all tANS codes in a batch are written, we write these offsets.
@@ 2.3.1 Bin Optimization
Since each additional bin incurs some metadata storage, it is usually suboptimal to store all histogram bins. When adjacent bins have similar probability density, they can be combined without increasing the binned data’s bit cost by much. We use a well-known $\mathcal{O}(k^{2})$ dynamic programming algorithm [21] to partition the bins optimally, minimizing bit cost. This algorithm requires only the bins and a cost function – in our case, the bit cost $\hat{H}_{i:j}$ associated with an optimized bin formed from unoptimized bins $i$ through $j$:
$\hat{H}_{i:j}=M+n_{i:j}\cdot(\alpha_{i:j}+\beta_{i:j})$
Here, $M$ is a constant for the number of bits required to store the bin’s metadata, $n_{i:j}=\sum_{l=i}^{j}n_{l}$ is the count of latents in the bin, $\alpha_{i:j}=\log_{2}\left(n/n_{i:j}\right)$ is the idealized average entropy coding bits per latent in the bin, and $\beta_{i:j}=\lceil\log_{2}(b_{j}-a_{i}+1)\rceil$ is the offset bits per latent in the bin. Essentially, the bin pays its metadata cost and, for each number in its range, entropy coding and offset cost. Once the optimal bin partitioning is chosen, we replace each partition’s bins $(a_{i},b_{i},n_{i}),\ldots,(a_{j},b_{j},n_{j})$ with a single bin
$(a^{\prime},b^{\prime},n^{\prime})=\left(a_{i},b_{j},\sum_{l=i}^{j}n_{l}\right)$
@@ 2.4 Format
Pco is most flexible within a wrapping format. We found it important to decouple Pco’s goal of good compression from the varied access patterns and logical transformations required by real-world applications, which may include nullability, page statistics, or additional indices. To support these needs, Pco’s API produces three components that can be interleaved as needed.
@@
Every chunk contains one or more pages, which contain the actual numerical data (Figure 2). Pages store initial states for tANS and delta decoding, so when paired with the correct chunk metadata, they contain all information necessary to decompress their contents. Therefore pages can be decompressed in any order; it is not necessary to decompress preceding pages. Each page can contain multiple batches, the unit of decompression. Batches consist of all latents for 256 numbers and exist solely for performance purposes. Batches within a page must be decompressed serially.
@@ 3.1 Theoretical Results
In Appendix A, we obtain a bound on binned bit cost per number:
Theorem 1
Suppose $X$ is a mixture over domain $\{0,\ldots,T-1\}$ of $s$ disjoint integer distributions, each of which has a monotonic PMF. Then for any $k>2s$, there exists a binning of at most $k$ bins such that the expected binned bit cost $\hat{H}$ of a random draw from $X$ satisfies
$\hat{H}\leq H+\frac{3s\log_{2}(T)}{k-2s}\frac{T}{T-1}$
where $H$ is the base-2 Shannon entropy of $X$.
Informally, we call $X$ smooth if $s$ is small. And this is the case for many types of data; geometric distributions have $s=1$, Poisson distributions have $s=2$, and float normal distributions have $s=4$ when represented as ordered, unsigned integer latents. Since the Shannon entropy is the theoretical minimum required average bit cost, this theorem demonstrates $\mathcal{O}(1/k)$ convergence to optimal compression on SIID data. We prove this theorem by construction, and the binning that attains this bound uses bins of approximately equal probability mass. This lines up with our approach of using a histogram to assign unoptimized bins. Note that for any reasonable data type, $T/(T-1)\approx 1$, so in practice the bound is simply
$\hat{H}\leq H+\frac{3s\log_{2}(T)}{k-2s}.$
Pco converges faster than this bound in reality (Figure 3). This is to be expected, since our proof uses antagonistic assumptions. However, this bound can still be practical. For instance, using the default compression level’s 256 bins on a 64-bit data type with $s=1$, we guarantee that binning wastes fewer than $0.76$ bits per latent on average.